• DNS in Tampa issue 4-1-2019

    - Tampa DNS servers experienced an issue resolving domains. We believe this issue is now resolved, please reach out to support if you are still experiencing any issues.

  • 3-28-19 Denvery Elynk Down -RESOLVED

    Hello, The Elynk is back online and the outages should be cleared. If you see any further issues please let us know by contacting your support team. Thank you.

  • 3-28-19 Denvery Elynk Down

    Hello, We have reports that the Elynk to our Denver DC is down. This is effecting multiple customers ability to use their direct connection to AWS. As well as other connection issues. We are actively engaging with Level3 to diagnose the issue.

  • Newark Datacenter Outage - 11/15/2018 (Update)

    We have determined the cause of the outage was due to a hardware failure. Everything is currently up and running. Replacement parts have been ordered and will be installed as soon as possible.

  • Newark Datacenter Outage - 11/15/2018

    We have determined there is a limited outage in the Newark Datacenter (NWK03). We are working with our internal teams to get more information now.

  • Phone issues 10-30-2018

    Hosting support is available via phones now. All phone service has been restored. Tickets and Chats are still available through the customer portal.We apologize for the inconvenience.

  • Phone system issues 10/30/2018

    Our phone provider is currently offline right now and phone calls into support are not working. If you need to contact support you can log into the customer portal and either open a chat or a ticket.

  • Phone System Issue Resolved 10/21/2018

    Phone System has been restored to working state. We apologize for the inconvenience.

  • Phone System Outage - 10/21/2018

    Our Phone System is currently down. If you require assistance, please contact support via ticket or chat through the customer portal at

    We are in contact with our Phone provider and are working on resolving this issue as quickly as possible.

  • Internal Monitoring Issue: 9-27-2018

    We have resolved the issue with monitoring and have resumed normal operations.